# Occupied Processes: Going with the Flow

### v4

@@ 5.2 Local Occupied Volatility (LOV) Model
for some path-dependent feature $F_{t}(\omega)=\textnormal{f}(t,\omega)\in\operatorname{\mathfrak{F}}$, $\textnormal{f}\mathrel{\mathop{\ordinarycolon}}[0,T]\times\Omega_{T}\to\operatorname{\mathfrak{F}}$. The feature space $\operatorname{\mathfrak{F}}$ may be finite-dimensional, e.g., $F_{t}\in\{\sup_{s\leq t}X_{s},\int_{0}^{t}X_{s}ds\}$ in Brunick and Shreve [20],or infinite-dimensional in occupied volatility models where $\operatorname{\mathfrak{F}}=\operatorname{\mathcal{M}}$; see Section 5.2.1. Recall the smile calibration condition
@@
with the Fréchet derivative $D\sigma^{2}(\cdot,X_{t})\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathfrak{F}}\to\operatorname{\mathfrak{F}}^{*}$ and dual pairing $\operatorname{\mathfrak{F}}^{*}\times\operatorname{\mathfrak{F}}\ni(\ell,F)\to\ell\cdot F$, provided that $\operatorname{\mathfrak{F}}$ is a linear space. The derivative $D\sigma^{2}(\hat{F}_{t},X_{t})$ can be seen as the sensitivity of variance with respect to non-Markovian shocks. Noting that $\sigma^{2}(\hat{F}_{t},X_{t})^{2}$, $D\sigma^{2}(\hat{F}_{t},X_{t})$ are functions of $(t,X_{t})$, we can postulate that
$\sigma^{2}(F_{t},X_{t})=\sigma_{\text{loc}}^{2}(t,X_{t})+\ell(t,X_{t})\cdot(F_{t}-\hat{F}_{t}),\quad\ell\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}_{+}^{2}\to\operatorname{\mathfrak{F}}^{*},$ (71)
by replacing $\sigma^{2}(\hat{F}_{t},X_{t})$ in (70) with the local variance. Since $\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[F_{t}-\hat{F}_{t}|X_{t}]=0$, the above specification automatically satisfies the smile calibration condition (69) for any reasonable choice of $\ell$, e.g., such that $\sigma^{2}(F_{t},X_{t})$ remains positive. The local path-dependent volatility model (71) is reminiscent of local stochastic volatility (LSV) models [54, 55] where one multiplies $\sigma_{\text{loc}}$ by an exogenous stochastic processes $\alpha_{t}$ and normalized such that (69) holds. The volatility process is thus of the form
@@
where the conditional expectation is typically estimated using the particle method of Guyon and Henry-Labordère [54]. It is a powerful family of models that is consistent with vanilla options and capable of producing more realistic spot-vol dynamics than the standard local volatility model. On the other hand, local path-dependent volatility models also preserve market completeness by explaining future volatility from the price history directly. Moreover, considering an additive correction term $\ell(t,X_{t})\cdot(F_{t}-\hat{F}_{t})$ term is arguably preferred numerically over the fraction in (72) with distribution-dependent denominator.
@@ 5.2.1 The LOV Model
Let us set $F_{t}=\operatorname{\mathcal{O}}^{\kappa}_{t}$. The feature space is $\operatorname{\mathfrak{F}}=\operatorname{\mathcal{M}}$, and the Fréchet differential in (70) becomes the linear derivative $\delta_{\mathfrak{o}}$ seen in Section 3.1. Recalling the dual pairing $\phi\cdot\mathfrak{o}=\int_{\operatorname{\mathbb{R}}}\phi(x)\mathfrak{o}(dx)$, $(\phi,\mathfrak{o})\in\operatorname{\mathcal{C}}_{b}(\operatorname{\mathbb{R}})\times\operatorname{\mathcal{M}}$, we can then rewrite (70) as
@@
Note that $\delta_{\mathfrak{o}}\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa},X_{t})(x)=\frac{d}{dh}\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa}+h\delta_{x},X_{t})|_{h=0}$ describes the change in instantaneous variance when the asset spends extra time at some level $x$. We then replace $\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa},X_{t})$ in (73) by the local variance and write $\delta_{\mathfrak{o}}\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa},X_{t})(x)=\gamma_{t}^{\kappa}\ell(t,X_{t},x)$, with the inverse total mass $\gamma_{t}^{\kappa}=\operatorname{\mathcal{O}}_{t}^{\kappa}(\operatorname{\mathbb{R}})^{-1}=\frac{\kappa}{e^{\kappa t}-1}$ if $\kappa>0$ and $\gamma_{t}^{0}=\frac{1}{t}$. This leads finally to the Local Occupied Volatility (LOV) model,
@@
$\sigma^{2}(\mathcal{O}_{t}^{\kappa},X_{t})$ $=\sigma_{\text{loc}}^{2}(t,X_{t})+\gamma_{t}^{\kappa}\int_{\operatorname{\mathbb{R}}_{+}}\ell(t,X_{t},x)(\mathcal{O}_{t}^{\kappa}-\hat{\mathcal{O}}^{\kappa}_{t})(dx)$ (75)
The factor $\gamma_{t}^{\kappa}$ is introduced so that the magnitude of $\ell$ remains constant over time, and allows to interpreted the linear term as an average sensitivity of instantaneous variance as the realized occupation measure $\mathcal{O}_{t}^{\kappa}$ deviates from the implied one $\hat{\mathcal{O}}_{t}^{\kappa}$. The LOV model may be further simplified by assuming that $\ell$ is time-independent and, aligned with our experiments with neural OSDEs in the next section, that it depends on $X_{t},x$ only through the ratio $x/X_{t}$; see Example 14.
The simulation of (74)-(75) is carried out through an extension of the scheme in Section 5.1, where the projected flow $\hat{\operatorname{\mathcal{O}}}^{\kappa}$ is estimated using the particle method [54]. Concretely, the projected occupation times $\hat{O}_{m}(\omega_{j})$ in Step III.2. is estimated by the Nadaraya-Watson estimator,
$\hat{O}^{\kappa}_{m}(\omega_{j})=\frac{\sum_{j^{\prime}}\psi_{j,j^{\prime}}O^{\kappa}_{m}(\omega_{j^{\prime}})}{\sum_{j^{\prime}}\psi_{j,j^{\prime}}},\quad\psi_{j,j^{\prime}}=\psi(X_{t_{n}}(\omega_{j})-X_{t_{n}}(\omega_{j^{\prime}})),$
where $\psi\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}}_{+}$ is a kernel. In our numerical experiments, we use the quartic kernel $\psi(\delta)=\frac{15}{16}(1-(\delta/h)^{2})^{2}/h$, and set the bandwidth $h$ following the guidelines in [55, Chapter 11].
@@
Consider a time-homogeneous local volatility function given by the Markov projection $\sigma_{\text{loc}}(x)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\sigma^{2}(\operatorname{\mathcal{O}}_{t}^{\kappa},X_{t})|X_{t}=x]^{1/2},\,t\approx 0.5,$ displayed in Fig. 11, where $\sigma$ is the calibrated occupied volatility from Section 5.3. Moreover, assume that $\ell(t,X_{t},x)$ is time-homogeneous and depends on the ratio $x/X_{t}$, aligned with our findings in Section 5.3.3 below. Concretely, we consider the (scaled) hyperbolic tangent of the relative distance $1-x/X_{t}$ to mimick the shape of the sensitivities in Fig. 12. The hyperbolic tangent also ensures that the variance remains positive function if scaled accordingly. Indeed, if $\underline{\sigma}$ is the minimum of the local volatility function (typically positive) and $|\ell|\leq c<\underline{\sigma}^{2}/2$, then
$\sigma^{2}(\mathcal{O}_{t}^{\kappa},X_{t})\geq\underline{\sigma}^{2}-\gamma_{t}^{\kappa}\int_{\operatorname{\mathbb{R}}}|\ell(x/X_{t})|(\mathcal{O}_{t}^{\kappa}+\hat{\mathcal{O}}^{\kappa}_{t})(dx)\geq\underline{\sigma}^{2}-2c>0.$
@@
In practice, one would tune the sensitivity function $\ell$ to capture stylized facts of volatility, or match the price of other instruments such as barrier or VIX options. Fig. 7 shows two sample paths of the asset and volatilities where we choose $\kappa=12$ (one month window), and $T=0.5$, aligned with the last option’s expiry from Section 5.3. The right panel illustrates that local volatility is typically floored (by $\underline{\sigma}$). This undesirable property is overcome with the local occupied volatility due to the linear path-dependent term. Fig. 8 displays the occupation measures and sensitivity function $x\mapsto\ell(x/X_{t})$ (red curve) at $t=T$, together with the local (occupied) volatility. As can be seen, the local volatility is adjusted up or down according to the difference between the realized and implied occupation measures $\operatorname{\mathcal{O}}_{t}^{\kappa},\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa}$.
Algorithm 1 (Local Occupied Volatility (LOV) Model, Simulation)
$T>0$, $N\in\operatorname{\mathbb{N}}$ (# time steps), $M$ (# occupation time bins), $J\in\operatorname{\mathbb{N}}$ (# simulations)

### v5

@@ 5.2 Local Occupied Volatility (LOV) Model
for some path-dependent feature $F_{t}(\omega)=\textnormal{f}(t,\omega)\in\operatorname{\mathfrak{F}}$, $\textnormal{f}\mathrel{\mathop{\ordinarycolon}}[0,T]\times\Omega_{T}\to\operatorname{\mathfrak{F}}$. The linear feature space $\operatorname{\mathfrak{F}}$ may be finite-dimensional, e.g., $F_{t}\in\{\sup_{s\leq t}X_{s},\int_{0}^{t}X_{s}ds\}$ in Brunick and Shreve [20], or infinite-dimensional in occupied volatility models where $\operatorname{\mathfrak{F}}=\operatorname{\mathcal{M}}$; see Section 5.2.1. Recall the smile calibration condition
@@
with the Fréchet derivative $D\sigma^{2}(\cdot,X_{t})\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathfrak{F}}\to\operatorname{\mathfrak{F}}^{*}$ and dual pairing $\operatorname{\mathfrak{F}}^{*}\times\operatorname{\mathfrak{F}}\ni(\ell,F)\to\ell\cdot F$. The derivative $D\sigma^{2}(\hat{F}_{t},X_{t})$ is the sensitivity of variance with respect to path-dependent (or non-Markovian) shocks. Noting that $\sigma^{2}(\hat{F}_{t},X_{t})$, $D\sigma^{2}(\hat{F}_{t},X_{t})$ are functions of $(t,X_{t})$ and in light of (69), we can postulate the following local path-dependent volatility (LPDV) model,
$\sigma^{2}(F_{t},X_{t})=\sigma_{\text{loc}}^{2}(t,X_{t})+\ell(t,X_{t})\cdot(F_{t}-\hat{F}_{t}),\qquad\ell\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}_{+}^{2}\to\operatorname{\mathfrak{F}}^{*}.$ (71)
In particular, the term $\sigma^{2}(\hat{F}_{t},X_{t})$ in (70) has been replaced with the local variance. Since $\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[F_{t}-\hat{F}_{t}|X_{t}]=0$, the above specification automatically satisfies the smile calibration condition (69) for any reasonable choice of $\ell$, e.g., such that $\sigma^{2}(F_{t},X_{t})$ remains positive.
The LPDV model (71) is reminiscent of local stochastic volatility (LSV) models [54, 55] where one multiplies $\sigma_{\text{loc}}$ by an exogenous stochastic processes $\alpha_{t}$ and normalized such that (69) holds. The volatility process is thus of the form
@@
where the conditional expectation is typically estimated using the particle method of Guyon and Henry-Labordère [54]. It is a powerful family of models that is consistent with vanilla options and capable of producing more realistic spot-vol dynamics than the standard local volatility model. On the other hand, local path-dependent volatility models also preserve market completeness by explaining future volatility from the price history directly. Moreover, considering an additive correction term $\ell(t,X_{t})\cdot(F_{t}-\hat{F}_{t})$ term is arguably preferred numerically over the fraction in (72) with distribution-dependent denominator. Indeed, any error introduced when computing $\hat{F}_{t}$ is controlled in a linear way.
@@ 5.2.1 The LOV Model
Let us set $F_{t}=\operatorname{\mathcal{O}}^{\kappa}_{t}$. The feature space is $\operatorname{\mathfrak{F}}=\operatorname{\mathcal{M}}$, and the Fréchet differential in (70) becomes the linear derivative $\delta_{\mathfrak{o}}$ from Section 3.1. Recalling the dual pairing $\phi\cdot\mathfrak{o}=\int_{\operatorname{\mathbb{R}}}\phi(x)\mathfrak{o}(dx)$, $(\phi,\mathfrak{o})\in\operatorname{\mathcal{C}}_{b}(\operatorname{\mathbb{R}})\times\operatorname{\mathcal{M}}$, we can then rewrite (70) as
@@
The occupation sensitivity $\delta_{\mathfrak{o}}\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa},X_{t})(x)=\frac{d}{dh}\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa}+h\delta_{x},X_{t})|_{h=0}$ describes the change in variance when the asset spends more time than expected (captured by $\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa}$) at level $x$. We then replace $\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa},X_{t})$ with the local variance, and rewrite the occupation sensitivity as
$\delta_{\mathfrak{o}}\sigma^{2}(\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa},X_{t})(x)=\gamma_{t}^{\kappa}\ell(t,X_{t},x),\qquad\gamma_{t}^{\kappa}=\operatorname{\mathcal{O}}_{t}^{\kappa}(\operatorname{\mathbb{R}})^{-1}=\frac{\kappa}{e^{\kappa t}-1},\;\,\kappa>0,\qquad\gamma_{t}^{0}=\frac{1}{t}.$
Altogether, we have derived the Local Occupied Volatility (LOV) model,
@@
$\sigma^{2}(\mathcal{O}_{t}^{\kappa},X_{t})$ $=\sigma_{\text{loc}}^{2}(t,X_{t})+\gamma_{t}^{\kappa}\int_{\operatorname{\mathbb{R}}_{+}}\ell(t,X_{t},x)(\mathcal{O}_{t}^{\kappa}-\hat{\mathcal{O}}^{\kappa}_{t})(dx).$ (75)
The factor $\gamma_{t}^{\kappa}$ is introduced so as to maintain a constant magnitude of $\ell$ over time, and allow to interpret the linear term as the average sensitivity of variance when the realized occupation measure $\mathcal{O}_{t}^{\kappa}$ deviates from the implied one ($\hat{\mathcal{O}}_{t}^{\kappa}$).
Figure 7: Local occupied volatility (yellow dot) compared with local volatility (red). The yellow curve is the local sensitivity function $x\mapsto\ell(t,X_{t},x)$, which is integrated against the spread $\operatorname{\mathcal{O}}^{\kappa}_{t}-\hat{\operatorname{\mathcal{O}}}^{\kappa}_{t}$ between the realized occupation measure (blue) and the expected one (purple).
Without careful specification of the sensitivity function ($\ell$), the LOV model may produce negative variances because the path-dependent term in (75) is centered around zero. However, mild conditions on $\ell$ can guarantee $\sigma^{2}$ remains positive, as shown next.
Proposition 5.
[…]
Proof.
Let $c(t,x^{\prime})=\sup_{x}|\ell(t,x^{\prime},x)|$. Then $c(t,x^{\prime})<\frac{1}{2}\sigma_{\textnormal{loc}}^{2}(t,x^{\prime})$ by assumption, hence
$\sigma^{2}(\mathcal{O}_{t}^{\kappa},X_{t})\geq\sigma_{\textnormal{loc}}^{2}(t,X_{t})-\gamma_{t}^{\kappa}\int_{\operatorname{\mathbb{R}}}|\ell(t,X_{t},x)|(\mathcal{O}_{t}^{\kappa}+\hat{\mathcal{O}}^{\kappa}_{t})(dx)\geq\sigma_{\textnormal{loc}}^{2}(t,X_{t})-2c(t,X_{t})>0.$
Under the above assumption on $\ell$, the sensitivity function can nearly double the local variance or push it down to nearly zero. The path-dependent term in (75) thus offers a wide, yet controlled, range to adjust the local variance while ensuring positivity. Alternatively, one can rewrite the occupied variance in a multiplicative way, namely
$\sigma^{2}(\mathcal{O}_{t}^{\kappa},X_{t})=\sigma_{\text{loc}}^{2}(t,X_{t})\Big(1+\gamma_{t}^{\kappa}\int_{\operatorname{\mathbb{R}}_{+}}\tilde{\ell}(t,X_{t},x)(\mathcal{O}_{t}^{\kappa}-\hat{\mathcal{O}}^{\kappa}_{t})(dx)\Big),$ (76)
where the condition $|\tilde{\ell}(t,X_{t},x)|<\frac{1}{2}$ ensures positivity of $\sigma^{2}$.
@@
(one-factor LOV) Consider the multiplicative version (76), with $\tilde{\ell}(t,X_{t},x)=\beta\mathds{1}_{A}(x)$, $A\in\operatorname{\mathcal{B}}(\operatorname{\mathbb{R}}_{+})$, and $|\beta|<1/2$. Then,
$\sigma^{2}(\mathcal{O}_{t}^{\kappa},X_{t})=\sigma_{\text{loc}}^{2}(t,X_{t})\Big(1+\beta\gamma_{t}^{\kappa}\big(\mathcal{O}_{t}^{\kappa}-\hat{\mathcal{O}}^{\kappa}_{t}\big)(A)\Big).$
In view of Proposition 5, $\sigma^{2}>0$. Recalling that $\gamma_{t}^{\kappa}=\operatorname{\mathcal{O}}^{\kappa}_{t}(\operatorname{\mathbb{R}}_{+})^{-1}=\hat{\operatorname{\mathcal{O}}}^{\kappa}_{t}(\operatorname{\mathbb{R}}_{+})^{-1}$, the local variance is thus rescaled according to the fraction of time $\gamma_{t}^{\kappa}\mathcal{O}_{t}^{\kappa}(A)$ spent by the asset in $A$ compared to what is expected. Here, the occupied volatility depends only on $t,X_{t}$, and the occupation time $\operatorname{\mathcal{O}}^{\kappa}_{t}(A)$. As $(\operatorname{\mathcal{O}}^{\kappa}(A),X)$ is a two-dimensional Markov process, trajectories from the LOV model can thus be effectively simulated under this simple specification. In Algorithm 1, this would correspond to choosing $M=1$ corridor.
Examples for $A$ include the upside corridor $A^{+}=[X_{0},\infty)$, where a positive parameter $\beta$ would be consistent with the leverage effect. Alternatively, one can consider a narrow corridor around the money, e.g., $A=[0.9X_{0},1.1X_{0}]$, where high values for $\mathcal{O}_{t}^{\kappa}(A)$ indicate some mean reversion of the recent path toward the initial value. This mean reversion would, in turn, translate into lower volatilities than $\sigma_{\text{loc}}$ by choosing $\beta<0$.
Example 15.
(EMA LOV) Suppose that $\ell(t,X_{t},x)=\beta\log x$ for some parameter $\beta\in\operatorname{\mathbb{R}}$. Then, using (67), we can rewrite the local occupied volatility (75) as
$\sigma^{2}(\mathcal{O}_{t},X_{t})$ $=\sigma_{\textnormal{loc}}^{2}(t,X_{t})+\beta(\text{EMA}_{t}-\widehat{\text{EMA}}_{t}),$
$\text{EMA}_{t}$ $=\frac{1}{1-e^{-\kappa t}}\int_{0}^{t}\log X_{u}\kappa e^{-\kappa(t-u)}du,\hskip 17.00024pt\;\widehat{\text{EMA}}_{t}=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\hskip 0.56905pt\text{EMA}_{t}\ |\ X_{t}\hskip 0.56905pt].$
The trend $\text{EMA}_{t}-\widehat{\text{EMA}}_{t}$ is similar to the first offset process of Hobson and Rogers [59], where the projected EMA is replaced by the log spot $\log X_{t}$. Note that choosing $\beta>0$ makes the path-dependent correction term consistent with the leverage effect.
Figure 8: Simulated price path and volatility in the Local Occupied Volatility (LOV) model, with sensitivity function given in Example 16.
Example 16.
@@
Consider a time-homogeneous local volatility function given by the Markov projection $\sigma_{\text{loc}}(x)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\sigma^{2}(\operatorname{\mathcal{O}}_{t}^{\kappa},X_{t})|X_{t}=x]^{1/2},\,t\approx 0.5,$ displayed in Fig. 11. The occupied volatility $\sigma$ is calibrated to options on the S&P 500 Index (SPX); see Section 5.3. Assume that $\ell(t,X_{t},\cdot)$ is time-homogeneous, and nondecreasing in $x$. For simplicity, consider the following (scaled) hyperbolic tangent of the moneyness $x/X_{t}$,
$\ell(t,X_{t},x)=\frac{1}{4}\underline{\sigma}^{2}\text{tanh}(\alpha x/X_{t}),\qquad\underline{\sigma}=\inf_{x}\sigma_{\text{loc}}^{2}(x)>0,\qquad\alpha>0.$
@@
According to Proposition 5, the occupied variance remains positive. Fig. 8 shows two sample paths of the asset and volatilities with $\kappa=12$ (one month window), and $T=0.5$, aligned with the last option’s expiry from Section 5.3. The right panel illustrates that local volatility is typically floored (by $\underline{\sigma}$). This undesirable property is overcome with the local occupied volatility due to its path-dependent correction term.
Fig. 7 displays the occupation measures and sensitivity function $x\mapsto\ell(t,X_{t},x)$ (yellow curve) at $t=T$, together with the local and occupied volatility. As can be seen, the local volatility is adjusted up or down according to the difference between the realized and implied occupation measures $\operatorname{\mathcal{O}}_{t}^{\kappa},\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa}$. In the left panel of Fig. 7, the recent path has spent more time above the spot than expected, namely $\operatorname{\mathcal{O}}_{t}^{\kappa}(dx)>\hat{\operatorname{\mathcal{O}}}_{t}^{\kappa}(dx)$ for almost every $x>X_{t}$, indicating a possibly large drawdown resulting in higher volatility. Since the sensitivity function is positive above the spot and negative otherwise, the occupied volatility in the LOV model will be greater than the local volatility, further capturing the leverage effect. The right panel of Fig. 7 shows the opposite situation, where the spot is close to the asset’s recent high, which typically translates into lower volatilities.
5.2.2 Simulation
We sample trajectories from (74)-(75) through an extension of the algorithm in Section 5.1. Specifically, the projected flow $\hat{\operatorname{\mathcal{O}}}^{\kappa}$ is estimated using the particle method [54]. Concretely, the projected occupation times $\hat{O}_{m}(\omega_{j})$ in Step III.2. is estimated by the Nadaraya-Watson estimator,
$\hat{O}^{\kappa}_{m}(\omega_{j})=\frac{\sum_{j^{\prime}}\psi_{j,j^{\prime}}O^{\kappa}_{m}(\omega_{j^{\prime}})}{\sum_{j^{\prime}}\psi_{j,j^{\prime}}},\qquad\psi_{j,j^{\prime}}=\psi(X_{t_{n}}(\omega_{j})-X_{t_{n}}(\omega_{j^{\prime}})),$
where $\psi\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}}_{+}$ is a kernel. In our numerical experiments, we use the quartic kernel $\psi(\delta)=\frac{15}{16}(1-(\delta/h)^{2})^{2}/h$, and set the bandwidth $h$ following the guidelines in [55, Chapter 11]. The complete simulation scheme is provided in Algorithm 1.
Algorithm 1 (LOV Model, Simulation)
$T>0$, $N\in\operatorname{\mathbb{N}}$ (# time steps), $M$ (# corridors), $J\in\operatorname{\mathbb{N}}$ (# simulations)